# thanks: ORCID: $0$-0003-2020-0803thanks: ORCID: $0$-0001-7285-0707thanks: ORCID: $0$-0003-0665-266X

### v1

@@ I Introduction
Consequently, there is a clear and pressing need for a suite of reliable diagnostics and validation techniques specifically tailored for SBI outputs. Several valuable approaches have already been developed to address this need. Simulation-Based Calibration (SBC) Cook et al. (2006); Talts et al. (2018); Hermans et al. (2021) provides a necessary condition for correctness by checking whether, on average, the ranks of the true parameters under the inferred posteriors are uniformly distributed over many simulated datasets. Probability-Probability (PP) plots are often used to assess this calibration. More recently, techniques like Tests of Accuracy with Random Points (TARP) Lemos et al. (2023) have been introduced to quantify the concentration and calibration of posteriors by measuring the coverage probability of credible regions. Standard Bayesian techniques, such as posterior predictive checks, can also be adapted. Broadly, this involves comparing simulations generated from the synthetic model to the observed data to check for model misspecification, model comparison, or goodness-of-fit (see Refs. Jeffrey and Wandelt (2024); Gessey-Jones and Handley (2023); Mancini et al. (2023); Cranmer et al. (2016); Dalmasso et al. (2020); Heinrich (2022); von Wietersheim-Kramsta et al. (2024); Cannon et al. (2022); Schmitt et al. (2024); Huang et al. (2023); Anau Montel et al. (2025); Wehenkel et al. (2024); Kelly et al. (2025); Pierre et al. (2025) for several approaches developed to address this general comparison/misspecification challenge). In general, while these methods are extremely valuable, they often assess average performance over many simulations (like SBC), or rely on visual inspection (like PP-plots and some PPCs), and may not fully characterise the uncertainty or potential biases associated with the posterior approximation for a single, specific observation.
@@
The main idea of this work is to perform quality checks by computing and comparing discrepancy measures between the posterior estimates produced by each member of the ensemble. In particular, we highlight how the Kullback-Leibler (KL) divergence Kullback and Leibler (1951) can be efficiently computed in an SBI context and provides a robust and interpretable quantity for comparison. This concept draws a clear analogy with diagnostic techniques used in Markov Chain Monte Carlo (MCMC) methods. In MCMC, multiple chains are often run in parallel from dispersed starting points. The convergence of these chains is assessed using diagnostics such as the Gelman-Rubin statistic ($\hat{R}$elman and Rubin (1992), which compares the variance within individual chains to the variance between chains. High $\hat{R}$ values track differences between the chains and indicate a lack of convergence, suggesting the samples do not yet represent the target posterior. Similarly, we propose below using the consistency (or lack thereof) across an ensemble of SBI estimators as one diagnostic for the reliability of the resulting posterior approximation.
We note that, of course, ensemble learning itself is a well-established paradigm in ML, where multiple individual models are trained and their predictions combined or compared to achieve better performance, robustness, and uncertainty quantification than any single model could provide (see, e.g., Sagi and Rokach (2018) for a review). Common techniques in this regard include bagging, boosting, and stacking. By training an ensemble of SBI posterior estimators, we can not only assess convergence and stability but also potentially obtain more robust posterior estimates by combining the ensemble members, for instance, through mixture models or averaging, and gain insights into the approximation uncertainty. Ensemble methods have been explored before in the SBI context. For example, in Ref. Alsing et al. (2019), ensembles of likelihood estimators were used in an active learning framework to optimally generate new simulations with the acquisition protocol aimed at reducing likelihood uncertainty across the ensemble. In addition, Ref. Hermans et al. (2021) explored the coverage properties of ensembles, and Ref. Cannon et al. (2022) demonstrated the behaviour of different SBI algorithms under various levels of model misspecification.
@@ II.1 Principles of Ensemble Learning
Beyond improving aggregate predictive accuracy, ensembles inherently provide a mechanism for assessing model stability and uncertainty. Disagreement among ensemble members trained under nominally similar conditions (for example, using the same algorithm and data budget but different random seeds or bootstrap samples) can denote high variance in the learning process or sensitivity to specific training instances Lakshminarayanan et al. (2017). Although ensemble techniques have seen application within SBI, primarily aimed at improving the accuracy or robustness of the resulting posterior approximation Alsing et al. (2019); Hermans et al. (2021); Yao et al. (2024), their systematic utilisation as a diagnostic tool to gauge the reliability and convergence of the inference procedure for a single specific observation remains less explored. Addressing this problem is the central motivation for the present work.
@@ II.3 The KL Divergence Matrix for Ensemble Assessment
A practical advantage, particularly relevant for SBI methods based on neural density estimators for the posterior (such as NPE) [^1] In particular, while generating samples is simple and efficient in the NPE case, it can be more costly for alternative approaches like NRE or NLE., is the feasibility of estimating these KL divergences using Monte Carlo sampling. Specifically, $K_{ij}$ can be approximated by drawing $N_{s}$ samples [^2] To be more precise, enough samples should be drawn such that the Monte Carlo error in Eq. (3) falls below a chosen threshold. $\theta^{(k)}$ from the distribution $q_{i}(\theta|x_{\rm{obs}})$ and averaging the log-ratio of the probability densities:
@@
All this estimation requires is the ability to both sample from each $q_{m}(\theta|x_{\rm{obs}})$ and evaluate the log-densities $\log q_{m}(\theta|x_{\rm{obs}})$, capabilities readily available in many modern SBI implementations.
@@
II.4 Interpreting the KL divergence
@@
To understand how to interpret the KL divergence, we can use a toy model setup of two one-dimensional, Gaussian distributions with different means and variances. For this case, it is simple to derive an analytical expression for the KL divergence,
@@
where $\mu_{1}$, $\mu_{2}$, $\sigma_{1}$, and $\sigma_{2}$ are the means and standard deviations of the two distributions, respectively.
However, to better understand the mapping between the KL divergence and $\sigma$-levels, let us proceed further by considering small perturbations between the two distributions. We parametrize this as $\sigma_{2}=\sigma_{1}(1+\epsilon)$ and $\mu_{2}=\mu_{1}+\gamma\sigma_{1}$ such that $\epsilon$, $\gamma\ll 1$ are perturbations in units of $\sigma_{1}$. With this parameterisation, we can obtain the perturbative result,
@@
where $\vec{\mu}_{1,2}$ and $\Sigma_{1,2}$ are the mean vectors and covariance matrices respectively. In this case, we can gain insight into the perturbative limit using a similar approach to the one-dimensional case. We can diagonalise the system about the eigenvector basis $U$ of $\Sigma_{1}$ by assuming $\Sigma_{1}=U\Lambda^{2}U^{\dagger}$ with $\Lambda=\mbox{diag}(\sigma_{1},...,\sigma_{n})$ and defining perturbations $\Sigma_{2}=U\Lambda^{\prime 2}U^{\dagger}$ and $U^{\dagger}(\vec{\mu}_{2}-\vec{\mu}_{1})=\Gamma\Lambda\vec{1}$ where $\Lambda^{\prime}=\Lambda(\mathbb{I}+\mathcal{E})$, $\mathcal{E}=\mbox{diag}(\epsilon_{1},...,\epsilon_{n})$, and $\Gamma=\mbox{diag}(\gamma_{1},...,\gamma_{n})$ Then, in the diagonal basis, we recover
@@
$\left.\hskip 5.69046pt+\frac{\delta P_{1}^{2}(\theta)+\delta P_{2}^{2}(\theta)-2\delta P_{1}(\theta)\delta P_{2}(\theta)}{P(\theta)}\right]\,\textrm{d}\theta.$
@@
Figure 2: KL matrix as a training diagnostic. Left: Mean of the KL matrix as a function of the number of training epochs for different training set sizes (corresponding to bottom legend). Middle: Value of the KL divergences in the ensemble (averaged over the last 30 epochs) at the end of the training as a function of training dataset size. In dashed grey, we show the expected $1/n_{\rm Train}$ behaviour before saturating the intrinsic variance floor. Right: Value of the loss function at the end of the training process as a function of the training dataset size.
@@ III KL divergence applied to ensemble training
In this section, we restrict ourselves to the case of Neural Posterior Estimation (NPE) and explore some applications of the KL matrix as a diagnostic tool for SBI. [^3] For simplicity, for all the examples considered in this work, we have only considered the upper triangular part of the K matrix. We expect the results to remain qualitatively identical if the lower triangular part of the K matrix is also included. In Section III.1, we begin by reviewing the main ideas behind NPE. Then, in Section III.2, we discuss the use of the KL matrix to assess the convergence among different networks in an ensemble. Finally, in Section III.3, we explore the usage of the KL matrix to identify model misspecification.
@@ III.2 Assessing ensemble convergence and estimating systematic errors
Figure 3: Evolution of the posteriors. True posterior distribution (black) compared to samples drawn from the different networks in the ensembles. The different panels (top left to bottom right) correspond to different training dataset sizes (colour code as in Fig. 2).
@@
The main results for this subsection are presented in Fig. 2 and Fig. 3. They demonstrate a number of core properties that justify our statement that the KL matrix can be used as a diagnostic tool for ensembles of posterior approximations.
Firstly, if we look at the left panel of Fig. 2, we observe the behaviour of the average KL divergence in the KL matrix as a function of the training epoch for a number of different training set sizes. Specifically, we see that for all training set sizes, the KL divergences initially decrease as the networks learn a better approximation to the true posterior. Eventually, once the information contained in the training dataset has been fully utilised and the networks have converged, an asymptotic value of the KL divergences is reached. The key insight here is that this asymptotic value can be used to directly estimate the systematic error associated with the specific training configuration (i.e., training set size, training dynamics, etc.) via the interpretation discussed in Section II.4.
If we then look at the middle panel of Fig. 2, we see a more detailed picture of the asymptotic values reached in the KL matrices for each experiment. These demonstrate empirically a number of key ideas pointed out in Section II. In particular, we note that for small training dataset sizes, we observe the predicted scaling of the KL divergences with $1/n_{\mathrm{Train}}$. We also observe a second regime, however, where the KL divergences cease to decrease and instead saturate. This indicates that there is a fundamental systematic variation in the training process itself (which is, of course, expected when using techniques like stochastic gradient descent or similar optimisation schemes). This saturation value can be used in a number of ways. First, it can be utilised in a comparison between different training strategies (e.g., learning rate scheduling, hyperparameter choices, etc.) where lower asymptotic values indicate more stable performance. Second, it can be used (as pointed out above) as a method to estimate the systematic error in the posterior reconstruction, which is directly associated with the dynamics of training neural networks. We emphasise that this information cannot be obtained when training only a single network, but must be bootstrapped via a trained ensemble.
@@
Finally, for reference, in Fig. 3, we show samples drawn from the three networks in the ensemble at the end of the training process, after conditioning to the same reference observation. The different panels correspond to different training dataset sizes, which are chosen to match Fig. 2. In all the panels, we show the true posterior in black. Once again, we can appreciate that, while all panels show qualitatively good agreement between the true posterior and the approximants, the KL matrix provides a quantitative measure to assess the accuracy and consistency of the approximation.

### v2

# thanks: ORCID: 0000-0003-2020-0803thanks: ORCID: 0000-0001-7285-0707thanks: ORCID: 0000-0003-0665-266X

@@ I Introduction
Consequently, there is a clear and pressing need for a suite of reliable diagnostics and validation techniques specifically tailored for SBI outputs. Several valuable approaches have already been developed to address this need. Simulation-Based Calibration (SBC) Cook et al. (2006); Talts et al. (2018); Hermans et al. (2021) provides a necessary condition for correctness by checking whether, on average, the ranks of the true parameters under the inferred posteriors are uniformly distributed over many simulated datasets. Probability-Probability (PP) plots are often used to assess this calibration. More recently, techniques like Tests of Accuracy with Random Points (TARP) Lemos et al. (2023) have been introduced to quantify the concentration and calibration of posteriors by measuring the coverage probability of credible regions. Standard Bayesian techniques, such as posterior predictive checks, can also be adapted. Broadly, this involves comparing simulations generated from the synthetic model to the observed data to check for model misspecification or goodness-of-fit. A variety of methods have been developed to address this challenge, including approaches that learn statistics to distinguish between data from different models Anau Montel et al. (2025); Cranmer et al. (2016); Dalmasso et al. (2020); Heinrich (2022); Huang et al. (2023); Wehenkel et al. (2024), adaptations of Bayesian model comparison and evidence estimation Jeffrey and Wandelt (2024); Gessey-Jones and Handley (2024); Mancini et al. (2023), deviations in different SBI algorithms Cannon et al. (2022); Schmitt et al. (2024); Kelly et al. (2025), and diagnostics tailored to specific scientific applications von Wietersheim-Kramsta et al. (2025); Pierre et al. (2025). In general, while these methods are extremely valuable, they often assess average performance over many simulations (like SBC), or rely on visual inspection (like PP-plots and some PPCs), and may not fully characterise the uncertainty or potential biases associated with the posterior approximation for a single, specific observed dataset $x_{\rm{obs}}$.
@@
The main idea of this work is to perform quality checks by computing and comparing discrepancy measures between the posterior estimates produced by each member of the ensemble. In particular, we highlight how the Kullback-Leibler (KL) divergence Kullback and Leibler (1951) can be efficiently computed in an SBI context and provides a robust and interpretable quantity for comparison. This concept draws a clear analogy with diagnostic techniques used in Markov Chain Monte Carlo (MCMC) methods. In MCMC, multiple chains are often run in parallel from dispersed starting points. The convergence of these chains is assessed using diagnostics such as the Gelman-Rubin statistic ($\hat{R}$Gelman and Rubin (1992), which compares the variance within individual chains to the variance between chains. High $\hat{R}$ values track differences between the chains and indicate a lack of convergence, suggesting the samples do not yet represent the target posterior. Similarly, we propose below using the consistency (or lack thereof) across an ensemble of SBI estimators as one diagnostic for the reliability of the resulting posterior approximation.
We note that, of course, ensemble learning itself is a well-established paradigm in ML, where multiple individual models are trained and their predictions combined or compared to achieve better performance, robustness, and uncertainty quantification than any single model could provide (see, e.g., Sagi and Rokach (2018) for a review). Common techniques in this regard include bagging, boosting, and stacking. By training an ensemble of SBI posterior estimators, we can not only assess convergence and stability but also potentially obtain more robust posterior estimates by combining the ensemble members, for instance, through mixture models or averaging, and gain insights into the approximation uncertainty. Ensemble methods have been explored before in the SBI context. For example, in Ref. Alsing et al. (2019), ensembles of likelihood estimators were used in an active learning framework to optimally generate new simulations with the acquisition protocol aimed at reducing likelihood uncertainty across the ensemble. In addition, Ref. Hermans et al. (2021) explored the coverage properties of ensembles, and Ref. Cannon et al. (2022) demonstrated the behaviour of different SBI algorithms under various levels of model misspecification. More closely related to the current work, Ref. Elsemüller et al. (2023) proposed the general principle of using deep ensemble disagreement to diagnose sensitivities arising from both insufficient training convergence and model misspecification. Our work provides a thorough validation and in-depth analysis of this diagnostic approach using a KL-divergence-based disagreement metric.
@@ II.1 Principles of Ensemble Learning
Beyond improving aggregate predictive accuracy, ensembles inherently provide a mechanism for assessing model stability and uncertainty. Disagreement among ensemble members trained under nominally similar conditions (for example, using the same algorithm and data budget but different random seeds or bootstrap samples) can denote high variance in the learning process or sensitivity to specific training instances Lakshminarayanan et al. (2017). Although ensemble techniques have seen application within SBI, primarily aimed at improving the accuracy or robustness of the resulting posterior approximation Alsing et al. (2019); Hermans et al. (2021); Yao et al. (2024), their systematic utilisation as a diagnostic tool to gauge the reliability and convergence of the inference procedure for a single specific observed dataset remains less explored. Addressing this problem is the central motivation for the present work.
@@ II.3 The KL Divergence Matrix for Ensemble Assessment
A practical advantage, particularly relevant for SBI methods based on neural density estimators for the posterior (such as NPE), is the feasibility of estimating these KL divergences using Monte Carlo sampling. Specifically, $K_{ij}$ can be approximated by drawing $N_{s}$ samples [^1] To be more precise, enough samples should be drawn such that the Monte Carlo error in Eq. (3) falls below a chosen threshold. $\theta^{(k)}$ from the distribution $q_{i}(\theta|x_{\rm{obs}})$ and averaging the log-ratio of the probability densities:
@@
All this estimation requires is the ability to both sample from each $q_{m}(\theta|x_{\rm{obs}})$ and evaluate the log-densities $\log q_{m}(\theta|x_{\rm{obs}})$, capabilities readily available in many modern SBI implementations. While generating samples is simple and efficient in the NPE case, it can be more costly for alternative approaches like NRE or NLE. However, this does not mean that estimating the components $K_{ij}$ is also computationally expensive. For NRE, each ensemble member learns the ratio $r_{i}(\theta,x)\approx p(x|\theta)/p(x)$, leading to a posterior estimate $q_{i}(\theta|x_{0})\propto r_{i}(\theta,x_{0})p(\theta)$. The log-posterior difference needed for the KL calculation in Eq. (3) simplifies to $\log r_{i}(\theta^{(k)},x_{\rm{obs}})-\log r_{j}(\theta^{(k)},x_{\rm{obs}})$, as the prior $p(\theta)$ cancels. As such, we see that the primary cost for NRE is generating samples from each ensemble member with stochastic sampling, after which computing the KL matrix only requires evaluating the learned ratio. For NLE, each member learns the likelihood $\tilde{q}_{i}(x|\theta)\approx p(x|\theta)$. This means that the posterior approximation is given by $q_{i}(\theta|x)=\tilde{q}_{i}(x|\theta)p(\theta)/\tilde{p}_{i}(x)$, where $\tilde{p}_{i}(x)$ is the approximate evidence of the data $x$ for each ensemble member $i$. In this case, the log-posterior difference involves both the log-likelihood difference and the log of the Bayes factor between members. To compute this in full, this requires both posterior samples from each estimator $\tilde{q}_{i}(x_{\rm{obs}}|\theta)$ and an estimate of the evidence for each member $\tilde{p}_{i}(x_{\rm{obs}})$. In general, this latter component may be challenging, however, it can be computed along with posterior sampling using techniques like nested sampling. Thus, while the KL framework is most efficiently applied in NPE, it remains conceptually and practically applicable to NRE and NLE, with the main cost being the stochastic sampling step already required to obtain posterior constraints.
@@
II.4 Interpreting KL divergence as systematic uncertainty
@@
A key aspect of our framework is interpreting the KL divergence as a lower bound on the systematic uncertainty of the posterior approximation. To understand what a given KL value implies about the disagreement between posteriors, it is helpful to consider the case of two one-dimensional, Gaussian distributions. For this case, it is simple to derive an analytical expression for the KL divergence,
@@
where $\mu_{1}$, $\mu_{2}$, $\sigma_{1}$, and $\sigma_{2}$ are the means and standard deviations of the two distributions, respectively. Whether a given level of KL divergence is "too large" depends entirely on the application context. If an analysis requires parameter constraints more precise than this systematic uncertainty, then such a KL value indicates that the ensemble has not achieved sufficient consistency. The translation given in this section is most accurate for approximately Gaussian posteriors; for highly non-Gaussian or multimodal distributions, the KL value serves as a general inconsistency metric which can be benchmarked across the ensemble.
To better understand the mapping between the KL divergence and $\sigma$-levels, let us proceed further by considering small perturbations between the two distributions. We parametrize this as $\sigma_{2}=\sigma_{1}(1+\epsilon)$ and $\mu_{2}=\mu_{1}+\gamma\sigma_{1}$ such that $\epsilon$, $\gamma\ll 1$ are perturbations in units of $\sigma_{1}$. With this parameterisation, we can obtain the perturbative result,
@@
where $\vec{\mu}_{1,2}$ and $\Sigma_{1,2}$ are the mean vectors and covariance matrices respectively. In this case, we can gain insight into the perturbative limit using a similar approach to the one-dimensional case. We can diagonalise the system about the eigenvector basis $U$ of $\Sigma_{1}$ by assuming $\Sigma_{1}=U\Lambda^{2}U^{\dagger}$ with $\Lambda=\mbox{diag}(\sigma_{1},...,\sigma_{n})$ and defining perturbations $\Sigma_{2}=U\Lambda^{\prime 2}U^{\dagger}$ and $U^{\dagger}(\vec{\mu}_{2}-\vec{\mu}_{1})=\Gamma\Lambda\vec{1}$ where $\Lambda^{\prime}=\Lambda(\mathbb{I}+\mathcal{E})$, $\mathcal{E}=\mbox{diag}(\epsilon_{1},...,\epsilon_{n})$, and $\Gamma=\mbox{diag}(\gamma_{1},...,\gamma_{n})$ . Then, in the diagonal basis, we recover
@@
$\hskip-28.45274pt\left.\hskip 5.69046pt+\frac{\delta P_{1}^{2}(\theta)+\delta P_{2}^{2}(\theta)-2\delta P_{1}(\theta)\delta P_{2}(\theta)}{P(\theta)}\right]\,\textrm{d}\theta.$
@@
Figure 2: KL matrix as a training diagnostic. Left: The mean pairwise KL divergence, monitored during the training of single ensembles. Each curve represents one ensemble trained with a specific dataset size (see legend), demonstrating how the KL matrix can be used to track convergence and determine when further training is unlikely to improve consistency. Middle: Value of the KL divergences in the ensemble (averaged over the last 30 epochs) at the end of the training as a function of training dataset size. In dashed grey, we show the expected $1/n_{\rm Train}$ behaviour before saturating the intrinsic variance floor. Right: Value of the loss function at the end of the training process as a function of the training dataset size.
@@ III KL divergence applied to ensemble training
In this section, we restrict ourselves to the case of Neural Posterior Estimation (NPE) and explore some applications of the KL matrix as a diagnostic tool for SBI. [^2] For simplicity, for all the examples considered in this work, we have only considered the upper triangular part of the K matrix, as $K_{ij}$ and $K_{ji}$ provide correlated information. The asymmetry ($K_{ij}\neq K_{ji}$) can provide additional diagnostic information about the relative shapes of the posteriors, but we do not explore it in this work as we found it to be a subdominant effect in our experiments. In Section III.1, we begin by reviewing the main ideas behind NPE. Then, in Section III.2, we discuss the use of the KL matrix to assess the convergence among different networks in an ensemble.
Finally, in Section III.3, we explore the usage of the KL matrix to identify model misspecification.
@@ III.2 Assessing ensemble convergence and estimating systematic errors
Figure 3: Evolution of the posteriors. True posterior distribution (black) compared to samples drawn from the different networks in the ensembles. The different panels (top left to bottom right) correspond to different training dataset sizes (colour code as in Fig. 2). Each panel shows the results from a single ensemble trained with a specific dataset size; multiple ensembles are only needed to compare how systematic uncertainties vary with training conditions, not to estimate the uncertainty for a single setup.
@@
The main results for this subsection are presented in Fig. 2 and Fig. 3. They demonstrate a number of core properties that justify our statement that the KL matrix can be used as a diagnostic tool for ensembles of posterior approximations. A key use case is monitoring the KL matrix for a single ensemble during training. As shown in the left panel of Fig. 2, when the KL divergence stabilises, we can conclude that further training will not improve ensemble consistency, and the asymptotic KL value provides an estimate of the systematic uncertainty for that training configuration. To estimate the systematic uncertainty for a single, fixed training setup, only one ensemble of multiple members is required.
Firstly, if we look at the left panel of Fig. 2, we observe the behaviour of the average KL divergence in the KL matrix as a function of the training epoch for a number of different training set sizes. Specifically, we see that for all training set sizes, the KL divergences initially decrease as the networks learn a better approximation to the true posterior. Eventually, once the information contained in the training dataset has been fully utilised and the networks have converged, an asymptotic value of the KL divergences is reached. The key insight here is that this asymptotic value can be used to directly estimate the systematic error associated with the specific training configuration (i.e., training set size, training dynamics, etc.) via the interpretation discussed in Section II.4. In our plots, for visual clarity, we show the average of the off-diagonal KL matrix elements. However, the primary diagnostic for assessing convergence or flagging outliers is the maximum pairwise KL divergence, $\max_{i\neq j}K_{ij}$, as this value provides the lower bound on the approximation error.
If we then look at the middle panel of Fig. 2, we see a more detailed picture of the asymptotic values reached in the KL matrices for each experiment. These demonstrate empirically a number of key ideas pointed out in Section II. This panel also demonstrates how the KL matrix can be used to diagnose the source of disagreement. We observe two distinct regimes: (i) a data-limited regime at small $n_{\mathrm{Train}}$, where the KL divergence decreases as $1/n_{\mathrm{Train}}$ (in line with Eq. (9)), indicating that disagreement is driven by insufficient training data; and (ii) an optimisation-limited regime at large $n_{\mathrm{Train}}$, where the KL divergence plateaus. This floor reflects intrinsic variance from the training dynamics (e.g., random initialisation, stochastic gradient descent, architectural instabilities), showing that simply adding more data will not improve consistency. This saturation value can be used to compare different training strategies (e.g. learning rate scheduling, hyperparameter choices, etc.), with lower values indicating more stable performance, and to estimate the systematic error associated with the training dynamics itself. We emphasise that this information cannot be obtained when training only a single network, but must be bootstrapped via a trained ensemble.
@@
Finally, for reference, in Fig. 3, we show samples drawn from the three networks in the ensemble at the end of the training process, after conditioning to the same reference observation. The different panels correspond to different training dataset sizes, which are chosen to match Fig. 2. In all the panels, we show the true posterior in black. Once again, we can appreciate that, while all panels show qualitatively good agreement between the true posterior and the approximants, the KL matrix provides a quantitative measure to assess the accuracy and consistency of the approximation. We demonstrate the convergence of both the average, and maximum/minimum KL divergences across the ensemble in the Appendix. While visual inspection of these 2D marginals is feasible, the KL matrix is particularly powerful in high-dimensional problems. The KL divergence captures the full multivariate disagreement in a single scalar per ensemble member pair, providing a quantitative and automated diagnostic that can be integrated into training pipelines and systematic uncertainty budgets.